Solve by factoring:
x^2-5x=6
Thanks for your help and time ;)

Answer by edjones(8007)

x^2-5x=6
x^2-5x-6=0
What factors of -6, with different signs, add up to -5? Ans: -6, +1
(x-6)(x+1)=0
x=6, x=-1
.
Ed